Four coins are tossed. Let H be the random variable representing the number of the heads occur. Find the values and construct the probability distribution of the random variable H.

1
Expert's answer
2022-03-01T19:03:59-0500

Sample space: {HHH, HHT, HTH, THH, HTT, THT, TTH, TTT}.

"P(H=0)=\frac{1}{8}."

"P(H=1)=\frac{3}{8}."

"P(H=2)=\frac{3}{8}."

"P(H=3)=\frac{1}{8}."